OpenStudy (anonymous):

Find a cubic function with the given zeros. You must show your work -2, 3, -3

OpenStudy (allank):

Hint: If x=a and x=b are the zeros of a function, then the function is (x-a)(x-b)

OpenStudy (anonymous):

im lost :(

OpenStudy (allank):

An example always helps. Take the function f(x)=x^2-5x+6 Can you calculate the zeros of that function?

OpenStudy (allank):

Still there @babydoll332 ?

OpenStudy (anonymous):

12,-10

OpenStudy (anonymous):

start with \[(x+2)(x-3)(x+3)\] and multiply this mess out to get your answer

OpenStudy (anonymous):

way would i multiply with?

OpenStudy (anonymous):

**what

OpenStudy (anonymous):

i mean to multiply these factors together

OpenStudy (allank):

f(x)=x^2-5x+6 You need to factorize it to get f(x) = (x-3)(x-2) Then the zeros become 3 and 2. Now, what @satellite73 is telling you is to do the reverse. Does that help?

OpenStudy (anonymous):

start by multiplying \((x+3)(x-3)\) and when you are done, multiply the result by \(x+2\)

OpenStudy (anonymous):

3x^2+4x-9

OpenStudy (anonymous):

no I'm so completely LOST

OpenStudy (allank):

Okay, do you understand why the function will be (x+2)(x-3)(x+3)?

OpenStudy (anonymous):

yes

OpenStudy (allank):

Wonderful. so you only have a problem with the multiplication?

OpenStudy (anonymous):

the anserw is this right 3x^2+4x-9 of (x+2)(x-3)(x+3)

OpenStudy (allank):

Lemme work it out here: (x+2)(x-3)(x+3) = (x^2-3x+2x-6)(x+3) = x^3+3x^2-3x^2-6x+2x^2+6x-6x-6

OpenStudy (anonymous):

ok then?

OpenStudy (allank):

So, put like terms together and simplify. x^3+3x^2-3x^2-6x+2x^2+6x-6x-6 x^3+2x^2-6

OpenStudy (anonymous):

not to butt in, but it would be somewhat simpler to multiply \((x+3)(x-3)\) first because you end up with only two terms

OpenStudy (anonymous):

\((x+3)(x-3)=x^2+3x-3x-9=x^2-9\)

OpenStudy (allank):

True, @satellite73 =)

OpenStudy (anonymous):

then multiply \((x^2-9)(x+2)=x^2+2x^2-9x-18\)

OpenStudy (anonymous):

no like terms to collect, and you are done at that step

OpenStudy (anonymous):

except i made a typo, it should be \[(x^2-9)(x+2)=x^3+2x^2-9x-18\]

OpenStudy (anonymous):

ohh ok i get it then way do i do

OpenStudy (anonymous):

?

OpenStudy (allank):

Choose the method you are more comfortable with. Both are okay. Just two ways of doing the same thing. :)

OpenStudy (anonymous):

so after i do that I'm done?

OpenStudy (anonymous):

?

OpenStudy (allank):

Yes

OpenStudy (anonymous):

thank u

OpenStudy (allank):

You're welcome